OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? a. 3, 4, 5 b. 28, 26, 12 c. 17, 14, 6 d. all of the above

OpenStudy (anonymous):

if i remember, the 2 smaller #'s should be greater than (idk if equal too) the larger number

OpenStudy (anonymous):

so d.

OpenStudy (shinalcantara):

for a right triangle you'll use pythagorean theorem \[c^2 = a^2 + b^2 \] plug in to 'a' and 'b' the shorter sides since 'c' is the hypotenuse which is the longest side. if it gives the value of c then it is a right triangle

OpenStudy (shinalcantara):

let's check for choice a: \[c = \sqrt{3^2 + 4^2}\] what's the value of c?

OpenStudy (anonymous):

5

OpenStudy (shinalcantara):

then it's a right triangle. let's check choice b \[c = \sqrt{12^2 + 26^2}\]

OpenStudy (anonymous):

28.6356......

OpenStudy (shinalcantara):

does it correspond to the third side?

OpenStudy (anonymous):

i don't think so unless u round it down, but im not sure

OpenStudy (shinalcantara):

no? then it's not a right triangle. let's have choice c \[c = \sqrt{6^2 + 14^2}\]

OpenStudy (anonymous):

15.2315...

OpenStudy (shinalcantara):

so what do you think is the answer to that question above?

OpenStudy (anonymous):

A

OpenStudy (shinalcantara):

=)

OpenStudy (anonymous):

Thanks